OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

solve for x (x1 + x2)/2 = x_mid (x + (-2))/2 = 2 (x - 2)/2 = 2 ... ... x = ??

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

6-2=4/2=2

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

Thanks for the help

jimthompson5910 (jim_thompson5910):

you're welcome